OpenStudy (anonymous):

Find the cube roots of 125(cos 288° + i sin 288°). cubed root(125*(cos(288)+ i sin (288)) cubed root(125)*cubed root(cos(288)+i sin(288)) 5*(cos(288)+ i sin(288))^1/3

OpenStudy (anonymous):

@jdoe0001

OpenStudy (helder_edwin):

given \[\large z=\rho(\cos\phi+i\sin\phi) \] then the n n-th roots of z are given by: \[\Large w_k=\sqrt[n]{\rho}\left[\cos\frac{\phi+2k\pi}{n}+i\sin\frac{\phi+2k\pi}{n} \right] \] with k=0,1,...,n-1

OpenStudy (anonymous):

i just need to know how to finish it

OpenStudy (anonymous):

cubed root(125*(cos(288)+ i sin (288)) cubed root(125)*cubed root(cos(288)+i sin(288)) 5*(cos(288)+ i sin(288))^1/3

OpenStudy (anonymous):

^^that is what i have so far

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

\[\Large w_0=\sqrt[3]{125}\left(\cos\frac{288}{3}+i\sin\frac{288}{3}\right) \]

OpenStudy (helder_edwin):

u were doing it just fine. u need to apply now DeMoivre's formula to finish and get what i got.

OpenStudy (helder_edwin):

but that is just the first of the three cubic roots of the given complex number

OpenStudy (helder_edwin):

did u get it?

OpenStudy (jdoe0001):

\(\large { \bf \sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 0)}}}{{\color{blue}{ 3}}}}\qquad \\\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 1)}}}{{\color{blue}{ 3}}}}\qquad \\\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 2)}}}{{\color{blue}{ 3}}}} }\)

OpenStudy (jdoe0001):

well sorta missing a number there =) \(\bf \large {\sqrt[{\color{blue}{ 3}}]{125}\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 0)}}}{{\color{blue}{ 3}}}} \\ \quad \\ \sqrt[{\color{blue}{ 3}}]{125}\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 1)}}}{{\color{blue}{ 3}}}} \\ \quad \\ \sqrt[{\color{blue}{ 3}}]{125} \sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (2\pi \cdot 2)}}}{{\color{blue}{ 3}}}} }\)

OpenStudy (helder_edwin):

though \(2\pi\) should be replaced by 360º, since u r working with degrees.

OpenStudy (helder_edwin):

i forgot to mention that. sorry.

OpenStudy (jdoe0001):

yeap =)

OpenStudy (jdoe0001):

\(\bf \large {\sqrt[{\color{blue}{ 3}}]{125}\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (360^o \cdot 0)}}}{{\color{blue}{ 3}}}} \\ \quad \\ \sqrt[{\color{blue}{ 3}}]{125}\sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (360^o \cdot 1)}}}{{\color{blue}{ 3}}}} \\ \quad \\ \sqrt[{\color{blue}{ 3}}]{125} \sqrt[{\color{blue}{ 3}}]{\cfrac{288^o+{\color{brown}{ (360^o \cdot 2)}}}{{\color{blue}{ 3}}}} }\)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

sorry i was eating dinner

OpenStudy (anonymous):

22.9, 30, 34.7...are those the cubed roots? @jdoe0001 @helder_edwin